11-8 Castle Company provides estimates for its uncollectible accounts. The allowance for uncollectible accounts had a credit balance of $17,430 at the beginning of 2018 and a $22,710 credit balance at the end of 2018 (after adjusting entries). If the direct write-off method had been used to account for uncollectible accounts (bad debt expense equals actual write-offs), the income statements for 2018 would have included bad debt expense of $17,400 and revenue of $2,500 from the collection of previously written off bad debts.

Required: Determine bad debt expense for 2018 according to the allowance method.

Bad debt expense for the year of 2018__________________

Mountain High Ice Cream Company transferred $73,000 of accounts receivable to the Prudential Bank. The transfer was made with recourse. Prudential remits 90% of the factored amount to Mountain High and retains 10% to cover sales returns and allowances. When the bank collects the receivables, it will remit to Mountain High the retained amount (which Mountain estimates has a fair value of $6,300). Mountain High anticipates a $4,300 recourse obligation. The bank charges a 3% fee (3% of $73,000), and requires that amount to be paid at the start of the factoring arrangement.

Required: Prepare the journal entry to record the transfer on the books of Mountain High assuming that the sale criteria are met.(If no entry is required for a transaction/event, type "No journal entry required" in the first account field.)

Journal Entry Worksheet

Record the transfer of accounts receivable.

Event General Journal Debit Credit

1
Answer #1

1. (i). Amount written off without recovery as per direct method= $(17400-2500) = $14900

(ii). Increase in allowance for uncollectible accounts= (Amt. in the end- Amt. in beginning)= $(22710-17430)= $5280

(iii). Total Bad debts as per allowance method= $(14900+5280)= $20180

Journal Entry as per Allowance Method:

a. Written off Entry

Allowance for Doubtful Debts A/c Dr. 17400

To Account's Receivable A/c 17400

b. Recovery Entry

Account's Receivable A/c Dr. 2500

To Allowance for Doubtful Debts A/c 2500

Cash A/c Dr. 2500

To Account's Receivable A/c 2500

c. Adjustment Entry

Bad Debts A/c Dr. 20180

To Allowance for Doubtful Debts A/c 20180

2. a. Mountain High Ice Cream Company received 90% of the cash paid to Prudential = 73000*90% = $65700

Sales Return and Allowance = (73000-65700) = $7300

So, Journal Entry will be:

Cash A/c Dr.    65700

Sales Return & Allowance A/c Dr. 7300

To Account Receivable A/c 73000

b. No entry will be recorded in the books of Mountain High unless and until the bank collects more than 90% of the Account Receivable.

c. Bank fees = 3%*73000 = $2190

Factoring Fees A/c Dr. 2190

To Cash A/c 2190